OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos x

OpenStudy (anonymous):

@e.mccormick can you help me plz

OpenStudy (anonymous):

Use identity... sin2x=2sinxcosx.... So youre answer is 2sinxcosx - cosx (unless if you want to factor it a little by factoring out a cosx... You get cosx(2sinx-1)... That doesnt really matter though) the better answer is the one on the top

OpenStudy (anonymous):

@Gerardo_cast23 the answer is not avalable plz come back

OpenStudy (anonymous):

Allrighty. hmm wierd. What are the answers given?

OpenStudy (anonymous):

2 sin x cos2x sin x cos x (2 sin x - 1) 2 sin x

OpenStudy (anonymous):

Then yeah its the third one.. Look at the bottom paragraph of the first answer/response I gave. Do you understand it?

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can you help with another

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

Verify the identity. cot (x minus pi divided by two) = -tan x

OpenStudy (anonymous):

and can you show work plz

OpenStudy (anonymous):

ok give me a little bit of minutes; i havent taken trig for a year now lol

OpenStudy (anonymous):

Here's a little cheat sheet of Trig identities... The asnwer can be seen under co-function identities but I bet youre teacher wants the answer written out in a different manner. Wait a little

OpenStudy (anonymous):

Sorry forgot to attach the link lol http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

so what now?

OpenStudy (anonymous):

so it tanx?

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

Yeah sorry I was working it out And well it says to verify the equation....Heres the work in the next response. I think its tanx but your problem says its -tanx.. Lets see...

OpenStudy (anonymous):

\[\cot(x-\pi/2)= -tanx\] \[cotx=\frac{ cosx }{ sinx }\] \[\cot(x-\pi/2) = \frac{ cosx-\pi/2 }{ sinx-\pi/2 }\] SO use identity \[\cos(a-B)=cosacosB+sinasinB\] and\[\sin(a-B)=sinacosB-cosasinB\] (IM still writing but heres half of it

OpenStudy (anonymous):

so a is x... and B is the pie/2 so now we plug in and solve...\[\frac{ cosxcos \pi /2+sinxsin \pi /2 }{ sinxcos \pi/2 -cosxsin \pi/2 }\] Now solve for the ones you know.... cos pie/2=0 and sin pie/2 is 1.... you should be left with -sinx/cosx... which is the same as -tanx

OpenStudy (anonymous):

whew that was a long one

OpenStudy (anonymous):

can you help me with another, your good

OpenStudy (anonymous):

@Gerardo_cast23 ?

OpenStudy (anonymous):

haha thank you. and it depends how long is it.. I should be dowmstairs eatinglunch right now>.< I just got out of my free period/calculus class and now im in lunch

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ill post and you decide ok?

OpenStudy (anonymous):

Verify the identity (1+ sinx)/(cos x) + (cos x)/(1+sin x) = 2 sec x?

OpenStudy (anonymous):

is it too long?

OpenStudy (anonymous):

yeah I think it is. Im very sorry; I honestly dont know where to start on this problem... you pretty much have to guess and check... Sorry. yeah im going to go to Lunch. To start on this problem I would start at the left hand side; 2secx... by saying that.. Hopefully this helps instead of me explaining it http://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.424507.html